2021-03-21

Observers: Steve Williams, Timo Kravtsov , Claudia Gutiérrez,

Support: Takashi Nagao

TAT: Tassilo Schweyer, Jesper Sollerman, Sheng Yang

EFOSC2 Calibrations: bias frames (EFOSC2), Flats (SOFI)

Weather:

Start: humidity 51%, wind 5 m/s, seeing

Middle: humidity 26%, wind 7.9 m/s, seeing 1.24

End: humidity 24%, wind 9 m/s, seeing 1.38

Comments:

01:15: We start the night with a problem in the mirror (something with the detectors). Because of this, image analysis was not possible. The problem was not solved but we could start the observations.

05:49: The problems with the actorsA (in the mirror) continue. We changed to EFOSC and spent some time trying to solve the problem (doing some focus).